# Laser: Stratified Selective Sampling for Instruction Tuning with Dedicated Scoring Strategy

### v2

@@ Abstract
**footnotetext: These authors contributed equally to this work.
@@ 1 Introduction
Large language models (LLMs) can perform a wide range of text-based tasks through chat interfaces. Their generalist abilities stem from an extensive post-training phase, during which they are optimized to generate useful responses to user queries (Sanh et al., 2022; Mishra et al., 2022; Wei et al., 2021; Ouyang et al., 2022). The choice of training data in this phase has a major impact on model performance (e.g. Zhou et al., 2023a; Xia et al., 2024; Liu et al., 2024b; Chen et al., 2024; Ge et al., 2024).
Prior work identifies three key properties of effective instruction-tuning (IT) data (Zhou et al., 2023a): i) difficulty/relevance, reflecting how much a query contributes to learning (Li et al., 2024b; Liu et al., 2024a, c; Zhao et al., 2024b); ii) quality, the usefulness and accuracy of responses (Zhao et al., 2024a; Chen et al., 2024; Liu et al., 2024c); and iii) diversity, the scope of within-domain variability and cross-domain coverage of the data (Ge et al., 2024; Lu et al., 2024). Although many approaches address one or more of these aspects, they often face limitations that make them difficult to apply in practice.
@@ 2 Related work
Prior work on IT data selection generally falls into two categories based on how they assess sample difficulty, quality, and diversity: external scoring and model-inherent criteria. External scoring methods rely on: (i) handcrafted features such as coherence, grammaticality, naturalness and understandability (Cao et al., 2024); (ii) heuristics based on length or formatting (Zhao et al., 2024a; Muennighoff et al., 2025); and (iii) scores derived from LLMs of varying scales (LLM-scorers). Notable approaches leveraging LLM-scorers are: AlpaGasus (Chen et al., 2024), which prompts ChatGPT to score each data point; InsTag (Lu et al., 2024), which uses ChatGPT to generate open-ended tags for deriving complexity/diversity; Deita (Liu et al., 2024c), which fine-tunes LLaMA-1-13B on ChatGPT-annotated complexity/quality labels as Deita-scorers; and CaR (Ge et al., 2024), which trains a 550M reward model to rank instruction-response pairs by quality.
@@ 3 Methods
Previous research has shown that i) instruction difficulty (see e.g. Li et al., 2024b; Liu et al., 2024a, c; Zhao et al., 2024b) ii) response quality (see e.g. Zhao et al., 2024a; Chen et al., 2024; Liu et al., 2024c) and iii) diversity/composition (see e.g. Ge et al., 2024; Lu et al., 2024) are crucial for effective data selection for LLM finetuning. We address these three aspects in the three-step Laser pipeline as illustrated in Figure 1:
@@ 3.1 Instruction Classification
Inspired by the use-case categories defined in Ouyang et al. (2022), we establish the following task categories $\mathcal{L}$ for samples in instruction tuning datasets: i) Math, from simple calculation to problems requiring multi-step reasoning; ii) Coding, code generation tasks or programming-related question answering; iii) Generation, textual generation tasks including roleplaying, summarizing and rewriting passages; iv) Reasoning, questions requiring deductive/logical reasoning; v) Brainstorming, information-seeking and recommendation questions that require inductive reasoning, including classification tasks; vi) Factual QA, factual questions with simple facts as answers; and vii) Extraction, tasks requiring structured/answer extraction from textual contexts. We let two human annotators classify 80 samples from the popular MT-Bench dataset (Zheng et al., 2023a) and achieve high inter-annotator agreement (Cohen’s Kappa = 0.8635), demonstrating the discriminability of our categories. We compare two different approaches to build a classification model: i) LLM annotator and ii) SetFit classifier.
@@ 3.2.1 Difficulty scorer
Previous research suggests that difficulty (sometimes referred to as complexity) matters for data selection (see e.g. Liu et al., 2024c; Cao et al., 2024; Zhao et al., 2024b; Muennighoff et al., 2025), with more challenging data generally resulting in better model performance. However, existing difficulty metrics either lack generality across domains (e.g. length of reasoning trace in response in Muennighoff et al., 2025, or the Deita complexity scorer as shown in Figure 10(b)) or are strongly influenced by spurious features (e.g. the widely used Deita-complexity is strongly biased towards long sequences; see Figure 11; Liu et al., 2024c).
@@
We evaluate the effectiveness and the validity of the difficulty scoring. For effectiveness, we sample 25k data points from $\mathcal{D}$ (as it is described in Section 4.1) either at random or following the difficulty scores. We then finetune Mistral-7B-v0.3 (Jiang et al., 2023) on the resulting datasets. Appendix A.2.3 shows how the difficulty scorer helps improve performance on benchmarks independently of their domain. For validity (i.e. whether we are really measuring some notion of difficulty), we score the CodeForces section of DeepMind’s code-contests dataset (Li et al., 2022) and find high correlations with its human difficulty scores (details in Appendix A.2.4).
@@ 3.2.2 Quality scorer
Prior work shows that sample selection strategies based on response length or quality, as judged by external models, lead to better instruction tuning datasets (see e.g. Zhao et al., 2024a; Chen et al., 2024; Liu et al., 2024c). However, these strategies underestimate the diverse problem types within instruction tuning data, which may require different evaluation criteria. For example, the quality of responses to math and coding problems is heavily dependent on solution correctness, while constrained generation requires evaluation of adhered constraints. In this work, we designate a dedicated quality scorer for each category defined in Section 3.1.
@@
For Reasoning, Factual QA and Extraction, we employ the Deita quality scorer [^2] https://huggingface.co/hkust-nlp/deita-quality-scorer ($q_{\mathrm{deita}}$, Liu et al., 2024c), a finetuned LLaMA-13B for quality assessment.
@@
For Math data, we assess the soundness of mathematical reasoning using a specialised process reward model (Qwen2.5-Math-PRM-7B; Zhang et al., 2025b). Process reward models (PRMs; Lightman et al., 2023; Uesato et al., 2022) are trained to verify steps in reasoning traces as they are common in mathematical reasoning. We find double linebreaks and—if no double linebreaks present—single linebreaks as a delimiter to be a good heuristic to separate reasoning steps. As a reasoning trace breaks with a single erroneous step, we aggregate scores by taking the minimal score out of all steps within each trace, as the $q_{\mathrm{math}}$ score.
@@
To evaluate this scoring method, we use a 1K-sample test set from LiveCodeBench (Jain et al., 2024a), containing coding problems and LLM-generated responses. [^3] https://huggingface.co/spaces/livecodebench/code_generation_samples Using Qwen/Qwen2.5-Coder-14B-Instruct as the reviewing model, our framework achieves 70% accuracy and a 0.412 Pearson correlation with binary correctness labels (see Appendix A.3.2 for details).
@@ 3.3 Sampling
Greedily choosing the highest-scoring samples often leads to redundancy in some domains and under-representation in others. Thus, maintaining diversity in the final dataset $\mathcal{D}^{\prime}$ is essential. Since diversity is a property of the dataset as a whole—not of individual samples—selection should consider the dataset globally (e.g., Ge et al., 2024), rather than relying on iterative, sample-by-sample strategies (e.g., Liu et al., 2024c; Bukharin et al., 2024). Diversity can be promoted top-down by balancing category proportions (see e.g. Grattafiori et al., 2024; Dong et al., 2024), or bottom-up by ensuring sufficient semantic dissimilarity among selected samples (e.g., Liu et al., 2024c; Ge et al., 2024; Lu et al., 2024).
@@
We evaluate the effectiveness of our stratified sampling and clustering as we did for the difficulty and quality scorer evaluation in Sections 3.2.1 and 3.2.2. In addition, we demonstrate the efficiency and robustness of Laser’s sampling technique, particularly in comparison with Deita (Liu et al., 2024c) and CaR (Ge et al., 2024). The respective results are resported in Appendix A.4.1.
@@ 4.1 Experimental Setup
We use the aggregation of all the listed datasets in Table 2 as the source dataset $\mathcal{D}$ in all experiments, if not specified otherwise.
@@
While most experiments we finetune Mistral-7B-v0.3, we also demonstrate generalisation across models of varying sizes and families: i) tiiuae/Falcon3-10B-Base, ii) meta-llama/Llama-3.1-8B, iii) mistralai/Mistral-7B-v0.3, iv) Qwen/Qwen2.5-3B and v) HuggingFaceTB/SmolLM2-1.7B.
@@
    deita (Liu et al., 2024c)—ranking data points according to complexity and quality scores ($o_{deita}$ and $q_{deita}$, resp.), and iteratively builds $D^{\prime}$ by adding dissimilar samples based on embedding distances.
@@ 4.2.1 Main results
For each selection strategy, we construct a subset $\mathcal{D}^{\prime}={d}_{i=1}^{m}\subset\mathcal{D}$ of size $m=100,000$, and subsequently fine‑tune each of the five base models on $\mathcal{D}^{\prime}$. Figure 2 shows the resulting category distributions across sampling strategies, including the baselines. Figure 3(a) reports normalized score gains over the untuned base models for all evaluation benchmarks and models. Laser achieves the highest overall performance, surpassing all baselines and even the models trained on the entire source set ($|\mathcal{D}|>707\text{k}$). We examine the robustness by breaking the aggregated results down by model size (Figure 3(b)). Laser yields the most consistent gains of all tested conditions for all five bases and outperforms the full‑data training in all but the smallest model, confirming that the proposed sampling approach transfers well across models.
@@
To better understand the contribution of each Laser component, we ablate sampling by ranking with only difficulty scores ($f$), quality scores ($q$), or their combination (pp)—without clustering or stratification—and compare these variants to the full pipeline. As shown in Figure 4, using $f$ or $q$ alone yields clear improvements over random sampling. However, their full potential is realized only when combined within the complete Laser pipeline. Notably, relying solely on $p$ for ranking provides only marginal gains, and even underperforms $f$- or $q$-based sampling individually, highlighting the critical role of stratified sampling and clustering in Laser.
Hence, all three properties (difficulty, quality and diversity) are necessary but not sufficient on their own for effective and robust data sampling. This is intuitive: highly challenging queries answered incorrectly provide little benefit, while accurate responses to trivial questions yield weak training signals. The insufficiency of using solely $f$ or $q$ is especially interesting when we consider their covariance: in our source data $\mathcal{D}$, quality and difficulty are generally negatively correlated, with high negative correlations in Coding (r=.38r=-.38) and Math (r=.43r=-.43), and smaller correlations in other categories. This explains the large performance gaps observed for Coding and Math in Figure 4. We observe a similar relationship for the source datasets $\mathcal{D}_{weak}$ and $\mathcal{D}_{strong}$ in the following section.
@@ 4.2.2 Additional robustness experiments
As different users may require instruction data sets of different sizes, a selection pipeline should improve sampling at every target scale. To verify this, we fix the base model to Mistral‑7B‑Base and constructs datasets $\mathcal{D}^{\prime}$ of {1k, 5k, 10k, 25k, 50k} items using Laser as well as all baselines. Figure 5 demonstrates that our method outperforms the alternatives across scales, except when $m=1,000$. Notably, with only $m=50,000$ (7% of the source data), it surpasses the performance of full-data fine-tuning. We provide benchmark-specific plots in Figure 24 (Appendix A.7), reported without normalizing score gain.
@@
Next, it is very probable that the average effectiveness of the source data $\mathcal{D}$ changes from scenario to scenario. We therefore test Laser with two shifted source distributions: $\mathcal{D}_{\text{strong}}$ consisting of the full Tülu v3 (Lambert et al., 2024) – a meticulously optimised dataset, which can be considered highly effective – and $\mathcal{D}_{\text{weak}}$ composed of multiple datasets that have previously proved to be less effective for instruction tuning (see Appendix A.5.2 for details).
@@
Ultimately, $\mathcal{D}$ might be very homogeneous (i.e. contain mostly a specific data type). To simulate this, we create 5 skewed source sets $\mathcal{D}_{\text{skewed}}$ by randomly selecting two categories $l_{1},l_{2}\in\mathcal{L}$ for each $\mathcal{D}_{\text{skewed}}$ and retaining only items with labels $l_{i}\in{l_{1},l_{2}}$, along with a small (4%) residue from all other categories, yielding a strongly biased data distribution. We show the proportions of each category for all 5 new source datasets in Figure 7(a).
From each $\mathcal{D}_{\text{skewed}}$, we sample 25k items with the same strategies as previously and fine‑tune Mistral‑7B‑Base on each of the resulting $\mathcal{D}_{\text{skewed}}^{\prime}$. We expected diversity-aware sampling strategies like Deita and Laser to mitigate bias by enforcing either semantic spread and explicit quotas. Indeed, the diversity-aware sampling strategies on average outperform the naive methods, with Laser having an edge over Deita, in terms of performance improvements in Figure 7(b) Upon closer inspection, we find that only Laser is reducing the introduced bias of categories while Deita exacerbates it (see details in Appendix A.5.2). Next to that, we observe a key limitation of Deita’s sampling method in this setup: Only adding dissimilar examples leads to quick saturation when source data is homogeneous and, thereby, hitting a ceiling of possible samples that can be sampled with Deita.

### v3

@@ Abstract
footnotetext: Accepted at Findings of the Association for Computational Linguistics: EMNLP 2025.**footnotetext: These authors contributed equally to this work.
@@ 1 Introduction
Large language models (LLMs) can perform a wide range of text-based tasks through chat interfaces. Their generalist abilities stem from an extensive post-training phase, during which they are optimized to generate useful responses to user queries (Sanh et al., 2022; Mishra et al., 2022; Wei et al., 2022; Ouyang et al., 2022). The choice of training data in this phase has a major impact on model performance (e.g. Zhou et al., 2023a; Xia et al., 2024; Liu et al., 2024b; Chen et al., 2024; Ge et al., 2024).
Prior work identifies three key properties of effective instruction-tuning (IT) data (Zhou et al., 2023a): i) difficulty/relevance, reflecting how much a query contributes to learning (Li et al., 2024b; Liu et al., 2024a, b; Zhao et al., 2024b); ii) quality, the usefulness and accuracy of responses (Zhao et al., 2024a; Chen et al., 2024; Liu et al., 2024b); and iii) diversity, the scope of within-domain variability and cross-domain coverage of the data (Ge et al., 2024; Lu et al., 2024). Although many approaches address one or more of these aspects, they often face limitations that make them difficult to apply in practice.
@@ 2 Related work
Prior work on IT data selection generally falls into two categories based on how they assess sample difficulty, quality, and diversity: external scoring and model-inherent criteria. External scoring methods rely on: (i) handcrafted features such as coherence, grammaticality, naturalness and understandability (Cao et al., 2024); (ii) heuristics based on length or formatting (Zhao et al., 2024a; Muennighoff et al., 2025); and (iii) scores derived from LLMs of varying scales (LLM-scorers). Notable approaches leveraging LLM-scorers are: AlpaGasus (Chen et al., 2024), which prompts ChatGPT to score each data point; InsTag (Lu et al., 2024), which uses ChatGPT to generate open-ended tags for deriving complexity/diversity; Deita (Liu et al., 2024b), which fine-tunes LLaMA-1-13B on ChatGPT-annotated complexity/quality labels as Deita-scorers; and CaR (Ge et al., 2024), which trains a 550M reward model to rank instruction-response pairs by quality.
@@ 3 Methods
Previous research has shown that i) instruction difficulty (see e.g. Li et al., 2024b; Liu et al., 2024a, b; Zhao et al., 2024b) ii) response quality (see e.g. Zhao et al., 2024a; Chen et al., 2024; Liu et al., 2024b) and iii) diversity/composition (see e.g. Ge et al., 2024; Lu et al., 2024) are crucial for effective data selection for LLM finetuning. We address these three aspects in the three-step Laser pipeline as illustrated in Figure 1:
@@ 3.1 Instruction Classification
Inspired by the use-case categories defined in Ouyang et al. (2022), we establish the following task categories $\mathcal{L}$ for samples in instruction tuning datasets: i) Math, from simple calculation to problems requiring multi-step reasoning; ii) Coding, code generation tasks or programming-related question answering; iii) Generation, textual generation tasks including roleplaying, summarizing and rewriting passages; iv) Reasoning, questions requiring deductive/logical reasoning; v) Brainstorming, information-seeking and recommendation questions that require inductive reasoning, including classification tasks; vi) Factual QA, factual questions with simple facts as answers; and vii) Extraction, tasks requiring structured/answer extraction from textual contexts. We let two human annotators classify 80 samples from the popular MT-Bench dataset (Zheng et al., 2023) and achieve high inter-annotator agreement (Cohen’s Kappa = 0.8635), demonstrating the discriminability of our categories. We compare two different approaches to build a classification model: i) LLM annotator and ii) SetFit classifier.
@@ 3.2.1 Difficulty scorer
Previous research suggests that difficulty (sometimes referred to as complexity) matters for data selection (see e.g. Liu et al., 2024b; Cao et al., 2024; Zhao et al., 2024b; Muennighoff et al., 2025), with more challenging data generally resulting in better model performance. However, existing difficulty metrics either lack generality across domains (e.g. length of reasoning trace in response in Muennighoff et al., 2025, or the Deita complexity scorer222hkust-nlp/deita-complexity-scorer as shown in Figure 10(b)) or are strongly influenced by spurious features (e.g. the widely used Deita-complexity is strongly biased towards long sequences; see Figure 11; Liu et al., 2024b).
@@
We evaluate the effectiveness and the validity of the difficulty scoring. For effectiveness, we sample 25k data points from $\mathcal{D}$ (as it is described in Section 4.1) either at random or following the difficulty scores. We then finetune Mistral-7B-v0.3 (Jiang et al., 2023) on the resulting datasets. Appendix A.2.3 shows how the difficulty scorer helps improve performance on benchmarks independently of their domain. For validity (i.e. whether we are really measuring some notion of difficulty), we score the CodeForces section of DeepMind’s code-contests dataset (Li et al., 2022) and find high correlations with its human-annotated difficulty scores (details in Appendix A.2.4).
@@ 3.2.2 Quality scorer
Prior work shows that sample selection strategies based on response length or quality, as judged by external models, lead to better instruction tuning datasets (see e.g. Zhao et al., 2024a; Chen et al., 2024; Liu et al., 2024b). However, these strategies underestimate the diverse problem types within instruction tuning data, which may require different evaluation criteria. For example, the quality of responses to math and coding problems is heavily dependent on solution correctness, while constrained generation requires evaluation of adhered constraints. In this work, we designate a dedicated quality scorer for each category $l\in\mathcal{L}$ defined in Section 3.1.
@@
For Reasoning, Factual QA and Extraction samples, we employ the Deita quality scorer333hkust-nlp/deita-quality-scorer (Liu et al., 2024b), a finetuned LLaMA-13B for quality assessment, which yields the $q_{\mathrm{deita}}$ score.
@@
For Math data, we assess the soundness of mathematical reasoning using a specialised process reward model (Qwen2.5-Math-PRM-7B; Zhang et al., 2025). Process reward models (PRMs; Lightman et al., 2024; Uesato et al., 2022) are trained to verify steps in reasoning traces as they are common in mathematical reasoning. We find double linebreaks and—if no double linebreaks present—single linebreaks as a delimiter to be a good heuristic to separate reasoning steps. As a reasoning trace breaks with a single erroneous step, we aggregate scores by taking the minimal score out of all steps within each trace, as the $q_{\mathrm{math}}$ score.
@@
To evaluate this scoring method, we use a 1K-sample test set from LiveCodeBench (Jain et al., 2024a), containing coding problems and LLM-generated responses.444livecodebench/code_generation_samples Using Qwen/Qwen2.5-Coder-14B-Instruct as the reviewing model, our framework achieves 70% accuracy and a 0.412 Pearson correlation with binary correctness labels (see Appendix A.3.2 for details).
@@ 3.3 Sampling
Greedily choosing the highest-scoring samples often leads to redundancy in some domains and under-representation in others. Thus, maintaining diversity in the final dataset $\mathcal{D}^{\prime}$ is essential. Since diversity is a property of the dataset as a whole—not of individual samples—selection should consider the dataset globally (e.g., Ge et al., 2024), rather than relying on iterative, sample-by-sample strategies (e.g., Liu et al., 2024b; Bukharin et al., 2024). Diversity can be promoted top-down by balancing category proportions (see e.g. Grattafiori et al., 2024; Dong et al., 2024), or bottom-up by ensuring sufficient semantic dissimilarity among selected samples (e.g., Liu et al., 2024b; Ge et al., 2024; Lu et al., 2024).
@@
We evaluate the effectiveness of our stratified sampling and clustering as we did for the difficulty and quality scorer evaluation in Sections 3.2.1 and 3.2.2. In addition, we demonstrate the efficiency and robustness of Laser’s sampling technique, particularly in comparison with Deita (Liu et al., 2024b) and CaR (Ge et al., 2024). The respective results are reported in Appendix A.4.1.
@@ 4.1 Experimental Setup
We use the aggregation of all listed datasets in Table 2 as the source dataset $\mathcal{D}$ in all experiments, if not specified otherwise.
@@
While in most experiments we finetune Mistral-7B-v0.3, we also demonstrate generalisation across models of varying sizes and families: i) tiiuae/Falcon3-10B-Base, ii) meta-llama/Llama-3.1-8B, iii) mistralai/Mistral-7B-v0.3, iv) Qwen/Qwen2.5-3B and v) HuggingFaceTB/SmolLM2-1.7B.
@@
    deita (Liu et al., 2024b)—ranking data points according to complexity and quality scores ($o_{deita}$ and $q_{deita}$, resp.), and iteratively builds $D^{\prime}$ by adding dissimilar samples based on embedding distances.
@@ 4.2.1 Main results
For each selection strategy, we construct a subset $\mathcal{D}^{\prime}={d}_{i=1}^{m}\subset\mathcal{D}$ of size $m=100,000$, and subsequently fine‑tune each of the five base models on $\mathcal{D}^{\prime}$. Figure 2 shows the resulting category distributions across sampling strategies, including the baselines. Figure 3(a) reports normalized score gains over the untuned base models for all evaluation benchmarks and models. Laser achieves the highest overall performance, surpassing all baselines and even the models trained on the entire source set ($|\mathcal{D}|>707\text{k}$). We examine the robustness by breaking the aggregated results down by model size (Figure 3(b)). Laser yields the most consistent gains of all tested conditions for all five bases and outperforms the random sampling in all but the smallest model, confirming that the proposed sampling approach transfers well across models.
@@
To better understand the contribution of each Laser component, we ablate sampling by ranking with only difficulty scores ($f$), quality scores ($q$), or their combination (pp)—without clustering or stratification—and compare these variants to the full pipeline. As shown in Figure 4, using $f$ or $q$ alone yields clear improvements over random sampling. However, their full potential is realized only when combined within the complete Laser pipeline. Notably, relying solely on $p$ for ranking provides only marginal gains, and even underperforms $f$- or $q$-based sampling individually, highlighting the importance of diversifying the sampling through stratification and clustering in Laser.
Hence, all three properties (difficulty, quality and diversity) are necessary but not sufficient for sampling data that consistently performs well. This is intuitive: highly challenging queries answered incorrectly provide little benefit, while accurate responses to trivial questions yield weak training signals. Another salient result is the high performance of $q$-sampling for Coding and Math. While the soundness of responses (i.e. $q$-scores) is arguably more important in these domains, we also find that $q$- and $f$-scores in our source data $\mathcal{D}$ are generally negatively correlated, with especially high negative correlations in Coding (r=.38r=-.38) and Math (r=.43r=-.43). Shifting emphasis onto the difficulty ($f$-scores and $p$-scores) might therefore reduce the response quality in these domains and explain the performance gaps observed in Figure 4. We find similar relationships between quality and difficulty for the source datasets $\mathcal{D}_{weak}$ and $\mathcal{D}_{strong}$ in the following section.
@@ 4.2.2 Additional robustness experiments
As different users may require instruction data sets of different sizes, a selection pipeline should improve sampling at every target scale. To verify this, we fix the base model to Mistral‑7B‑Base and constructs datasets $\mathcal{D}^{\prime}$ of {1k, 5k, 10k, 25k, 50k, 100k} items using Laser, as well as random and all data as baselines. Figure 5 demonstrates that our method outperforms the alternatives across scales, except when $m=1,000$. Notably, with only $m=100,000$ (14% of the source data), it surpasses the performance of full-data fine-tuning. We provide benchmark-specific plots in Figure 24 (Appendix A.7), reported without normalizing score gain.
@@
Next, it is very probable that the average effectiveness of the source data $\mathcal{D}$ changes from scenario to scenario. We therefore test Laser with two shifted source distributions: $\mathcal{D}_{\text{strong}}$ consisting of the full Tülu v3 (Lambert et al., 2024)—a meticulously optimised dataset, which can be considered highly effective—and $\mathcal{D}_{\text{weak}}$ composed of multiple datasets that have previously proved to be less effective for instruction tuning (see Appendix A.5.2 for details).
@@
Ultimately, $\mathcal{D}$ might be very homogeneous (i.e. contains mostly a specific data type). To simulate this, we create five skewed source sets $\mathcal{D}_{\text{skewed}}$ by randomly selecting two categories $l_{1},l_{2}\in\mathcal{L}$ for each $\mathcal{D}_{\text{skewed}}$ and retaining only items with labels $l_{i}\in{l_{1},l_{2}}$, along with a small (4%) residue from all other categories, yielding a strongly biased data distribution. We show the proportions of each category for all five new source datasets in Figure 7(a).
From each $\mathcal{D}_{\text{skewed}}$, we sample 25k items with the same strategies as previously and fine‑tune Mistral‑7B‑Base on each of the resulting $\mathcal{D}_{\text{skewed}}^{\prime}$. We expected diversity-aware sampling strategies like Deita and Laser to mitigate bias by enforcing either semantic spread and explicit quotas. Indeed, the diversity-aware sampling strategies on average outperform the naive methods, with Laser having an edge over Deita, in terms of performance improvements as illustrated in Figure 7(b). Upon closer inspection, we find that only Laser reduces the introduced bias of categories whereas Deita exacerbates it (see Appendix A.5.2 for details). Next to that, we observe a key limitation of Deita’s sampling method in this setup: only adding dissimilar examples leads to quick saturation when source data is homogeneous and, thereby, hitting a ceiling of possible samples that can be sampled with Deita.